OpenStudy (anonymous):

Use the properties of exponents to simplify the expression. Write the answer in scientific notation. (5x10^6)(6x10^-5) all divided by (4x10^8)

OpenStudy (anonymous):

www.math.com/tables/algebra/exponents.htm try take a look of the site. u may learn more about the exponent properties.

OpenStudy (anonymous):

i dont understand how to do this problem

OpenStudy (anonymous):

first: if u have (a*x^b)(c*x^d), this equal to (a*c)x^(b+d)

OpenStudy (anonymous):

so, consider the top part first, change x=10,a=5,b=6,c=6,d=-5 u got me so far?

OpenStudy (anonymous):

even though its scientific notation i can still do it like this?

OpenStudy (anonymous):

scientific notation like what?

OpenStudy (anonymous):

well its all written in scientific notation. but if i just see it like an exponent problem i can do it like this right?

OpenStudy (anonymous):

yes. i used a,b,c,d,x as arbitrary. thats in general. u just have to subtitute them with numbers according to the questions.

OpenStudy (anonymous):

okay so i did all that. then what should i do?

OpenStudy (anonymous):

just to make sure, ur question is \[(5\times 10^{6})(6 \times 10^{-5}) / (4 \times 10^{8})\] am i right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

for upper part: u wil have 30*10. but 30=3*10 therefore, 30*10= 3*10*10= 3*10^2 then, use this: if u have (a*x^b)/(c*x^d), this equal to (a/c)x^(b-d) for (3*10^2)/(4*10^8)

OpenStudy (anonymous):

the answer should be 7.5 * 10^(-7)

OpenStudy (anonymous):

how -7? isnt it 2-8?

OpenStudy (anonymous):

because 3/4=0.75= 7.5 * 10^(-1)

OpenStudy (anonymous):

ohhh i get it. thank you so much!

OpenStudy (anonymous):

no problem. im not simply give u the answer straight away so u can learn how to do exponent problems on your own.